Yoda pushes a 23.5 kg chair across the floor with a for F = 17.4N [W]. If the chair moves at constant velocity, what is the coefficient of friction,μ between the chair and the floor ?


1
Expert's answer
2021-05-23T16:36:20-0400

First, write Newton's second law: the net force is zero (because there is no acceleration), but two forces are acting on the chair: the force of Yoda's attraction "F" and the force of friction "f":


"0=F-f,\\
f=\mu N,\\
N=mg.\\\space\\
f=\mu mg,\\
0=F-\mu mg,\\\space\\
\mu=\frac{F}{mg}=\frac{17.4}{23.5·9.8}=0.076."